Problem 3.1. Cardiac pacemakers contain electrical connections that are platinum pins soldered onto a substrate. The question of interest is whether different operators produce solder joints with the same strength. Twelve substrates are randomly assigned to four operators. Each operator solders four pins on each substrate, and then these solder joints are assessed by measuring shear strength of the pins Strength (lbs) Substrate 2 5.60 6.80 8.32 8.707.64 7.44 7.48 7.80 7.72 8.40 6.98 8.00 5.04 7.38 5.56 6.96 8.30 6.86 5.62 7.22 5.72 6.40 7.54 7.50 8.36 7.04 6.92 8.186.20 6.10 2.75 8.14 9.00 8.64 6.60 8.18 8.30 8.54 7.68 8.92 8.46 7.38 8.08 8.128.68 8.24 8.09 8.06 Operator Substrate 1 Substrate 3 4 a) Describe the design and write the model equation in both the means and effect:s parameterization. What dependent variable did you use? Define model equation terms State any assumptions about the random effects terms in the model equation Use PROC GLM to obtain the analysis of variance (ANOVA) table for these data. Test the null hypothesis that the mean solder joint strength is the same for the different operators. Express your conclusion to the hypothesis test in terms of the application (as opposed to a purely statistical conclusion) b) c)
Answer #1

Data frame of the randomized block design to study the joining strengths of the 4 operators: Operators A B C D E FGHIJK L 5.6Total Variation in the Randomized Block Design c-1 T-1 GM:- j-0 i- i,j - 7.409 Among-Group Variation in the Randomized Block Design C- r. Σ (X.-GM-167979 に0 SSA: ·一 一 Among-Block Variation in the Randomized Block Design c-1 [7.573] Σ (Xi,j) |6.6751 1 7.176 I 8.213 i-o Random Variation in the Randomized Block Design SSE-SST_ SSA-SSBL=35.709Degrees of Freedom: The Mean Squares in the Randomized Block Design MSA SSA Randomized Block FSTAT Statistic SSE 10821 SSBL = 5.063 dJBL 1.5271 MSBL- dJSE MSA1.411 MSE stat FSTAT Statistic for Block Effects MSBL_-4.679 MSE statAnalysis-of-Variance Table for the Randomized Block Desigin Source Degrees of freedonm Sum of squaresS Mean Squares (Variance) Among gropus(A) MSA=1.411 MSE df,-11 SSA 16.798 MSA 1.527 stat Among Blocks(BL) dfвґ3 SSBL-15.189 MSBL-5.063 MSBL_-4.679 statMSE Error dfse-33 SSE-35.709MSE-1.082 Total dfr=47 SST-67.695 ANOVA model i-group 4 j-factor 12 Yi GM+SSA +SSBL Y 7.409+16.798+15.189Testing for factor and interaction effects: Null hypothesis: All four operators produce the same joining strength. Alternative hypothesis: H, Not all μ are equal Atleast of them produced a different joining strength. F test for differences among the substrates: MSA-1.4113 Fstat MSE Fais the upper tail critical value from an F distribution with dfA-11 of freedom. and dfsE-33 degrees α: 0.05 Upper-tail critical value of F distribution: FqF (1-a),dfA,dfsE)-2.0933 P= 1-pF (Fstat, d.,dfse) =0.2141 We do not reject the null hypothesis at α:-0.05, level of significance. Because Fstat-1.4113 <F,-2.093 and P>α You conclude that there is no evidence of a difference in the mean strengths among the different substrates.F test for a difference among the operaters MSBLA.6789 stat MSE α:= 0.05 F,is the upper tail critical value from an F distribution with r-1 and re(n-1) degrees of freedom Upper-tail critical value of F distribution: Fa qF((1-a),dfBL,fE-2.8916 P-1-pF (Fstat , dfBL, dfse) = 0.0079 We reject the null hypothesis at a 0.05, level of significance. Because Fstat 4.6789> F-2.892 and P<α You conclude that there is an evidence of a difference among the operators and so the strength of the joints is not same for all the opertors, atleast there is one operator among the four operators doing the joints with a different joining strength